OpenStudy (anonymous):

what is the answer for the square root of xcubed - xsquared plus 2x -2=0

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

\[x \sqrt{x}-x^2+2x-2=0\] \[x(\sqrt{x}-x)+2(x-1)=0\] I don't know what to do from here.

OpenStudy (anonymous):

multiply the whole equation by x^2 and solve further u will get something like x =0,1,2

OpenStudy (anonymous):

so like in polynomial equation form as in rational root form too

OpenStudy (anonymous):

i guess it should be applicable in both conditions but i did in polynomial form. u multiply whole equation by x^2 so that the suare root of the x^3 is removed . then u have equation as : x^3 -x^4 + 2x^3 -2x^2=0 then , -x^4 +3x^3 -2x^2=0 taking -x^2 as common -x^2(x^2-3x+2)=0 or x^2(x^2-3x+2)=0 this leads to x^2=0 or x^2-3x+2=0 x^2=0 gives x=0 and x^2-3x+2=0 gives x=1,2 so this finally gives us x=0,1,2 hope fully its correct ..

OpenStudy (anonymous):

r u there